Question 4 4. Determine what the guess solution yp would be for the method of undetermined Coefficients, for each non homogeneous linear
Answer #1

_92 - sinat + t² ? O sin2t++31 Auxico y equestrian =0 y tuy Lor (D²+424 m2+4= >m= Izi Yalt) = , Coszt + $ Sinat Now since sin416) Elz 14 e 4 yil uylty +27 40² 40+ =lue oth Auxilary zequating y oum2 -4m+Leo m= 4 I/16-16 x 4 8 Yect) = Get Cat e is comm